OpenStudy (anonymous):

Without finding the solution, how many solution do the following system has? Justify your solution. a. 0 = x^2 + 3x -3y -5 0 = 3x - 2y +2 b. y = x^2 + 2x -1 y= 2x -1 help i don't know how to do this...

OpenStudy (tkhunny):

Line vs. Parabola 0, 1, or 2 No other choices.

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

 The drawing shows you a parabola and three different lines. They will equal each other wher they intersect.

OpenStudy (tkhunny):

You could substitute or subtract. b. y = x^2 + 2x -1 y= 2x -1 Leads to 0 = x^2 How many Real Solutions are there to that?